OpenStudy (anonymous):

I have a work sheet from my algebra 2 class involving greater than and less than signs and im confused can anyone help?

OpenStudy (anonymous):

\[-12\le d +7\]

OpenStudy (marihelenh):

You need to isolate the variable. So, subtract 7 from both sides.

OpenStudy (anonymous):

What do you get

OpenStudy (anonymous):

i forgot the steps and the special tricks that are specificaly used when it involves a greater than or less than signs @twistnflip

OpenStudy (marihelenh):

The only thing you really have to remember is that when multiplying or dividing by a negative number, the sign flips.

OpenStudy (anonymous):

\[-19\le d\]

OpenStudy (marihelenh):

Yes, that is correct! You don't have to do any graphing or such right?

OpenStudy (anonymous):

nope! Thank you i have a couple more if i can give you my final answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

go ahead I will check them

OpenStudy (anonymous):

okay i gotta due them ill tag you guys when im done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[1z-1\ge 2z-3\]

OpenStudy (anonymous):

and i got

OpenStudy (anonymous):

\[z \ge 2\]

OpenStudy (anonymous):

eek, check your work on that one, I got something different

OpenStudy (marihelenh):

What was your work @mmrachel ?

OpenStudy (anonymous):

JK nevermind, I need to check my work. You are correct, sorry about that

OpenStudy (anonymous):

so i divided the negative 1 and didnt change any of the signs was i supposed to?

OpenStudy (marihelenh):

If you divide by a negative, you flip the sign.

OpenStudy (anonymous):

correct can you please type your work, it may be easier to check it

OpenStudy (anonymous):

ok i will

OpenStudy (marihelenh):

You had the correct answer, the sign just needed to be flipped.

OpenStudy (anonymous):

so instead of greater than its less than?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{-2 }{ 3 }r \le \frac{ 7 }{ 12 }\]

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

\[r \le \frac{ 21 }{ 24 }\]

OpenStudy (anonymous):

eek ok, I actually check my work on this one and that is wrong, mind typing your steps?

OpenStudy (anonymous):

i times both sides by the recipricole 3/-2 on both sides then i just multiplied 7/12 by 3/-2

OpenStudy (anonymous):

ok well silly, you forgot to simplify lol. what do 21 and 24 have in common

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

it simplifies to -7/8 (you accidentally cancelled the negative) final answer would be \[r \ge \frac{ -7 }{ 8 }\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yeah! lol i didnt do half bad. thanks again,

OpenStudy (anonymous):

no problem!!

OpenStudy (anonymous):

that was it

OpenStudy (anonymous):

great where should I post my english 2? I will post it in english and tag you will you come try?

OpenStudy (anonymous):

yes i will